OpenStudy (anonymous):

evaluate sin 3pi/4 in terms of pi (will award medal if you help me out)

OpenStudy (anonymous):

this is what you do locate \(\frac{3\pi}{4}\) on the unit circle on the last page of the attached cheat sheet then find the coordinates of the point corresponding to that angle the first coordinate is cosine and the second coordinate is sine

OpenStudy (anonymous):

let me know what you get btw the "in terms of \(\pi\)" makes no sense at all

OpenStudy (anonymous):

okay it says \[(\frac{-\sqrt2}{2}, \frac{ \sqrt2 }{ 2 } )\]

OpenStudy (anonymous):

and the "in terms of pie" was on my worksheet thing

OpenStudy (anonymous):

yes it certainly does so \[\sin(\frac{3\pi}{4})=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

really? can you post a screen shot? i would love to see it

OpenStudy (anonymous):

oh thats it?

OpenStudy (anonymous):

yup and \[\cos(\frac{3\pi}{4})=-\frac{\sqrt2}{2}\] and also \[\tan(\frac{3\pi}{4})=-1\] as it is sine over cosine easy right?

OpenStudy (anonymous):

all you need to do is find the coordinate of the point corresponding to the angle, and remember that the first coordinate is cosine and the second is sine i still want to know what "in terms of \(pi\)" means is your teacher a bit confused?

OpenStudy (anonymous):

i'm not sure... its an online course. i have no face to face contact with whatever makes the worksheets

ganeshie8 (ganeshie8):

i think they want to use pi as base unit

ganeshie8 (ganeshie8):

just scale ur answer by pi

OpenStudy (anonymous):

let \(f(x)=x^2+x\) find \(f(3)\) in terms of \(\pi\)

OpenStudy (anonymous):

makes no damned sense at all sine is a function and \[\sin(\frac{3\pi}{4})=\frac{\sqrt2}{2}\] a number

OpenStudy (anonymous):

it is not "in terms" of anything

OpenStudy (anonymous):

let me guess FLVS idiots strike again?

OpenStudy (freckles):

\[\sin(\frac{3\pi}{4})=\frac{\sqrt{2}}{2}+0 \pi \text{ lol }\]

ganeshie8 (ganeshie8):

some teachers are too smart they want students to compare the value of sin in terms of pi because pi is a measure of half the circumference of unit circle

OpenStudy (anonymous):

sine is a function put in a number get out a number

OpenStudy (anonymous):

@giniasal can you PLEASE post a screen shot? i would love to see the question and the question id number

ganeshie8 (ganeshie8):

\[\dfrac{\sqrt{2}}{2 } \approx 0.225 \pi \]

OpenStudy (anonymous):

lord

OpenStudy (anonymous):

it is instructed before the "practice" begins that we do everything "in terms of pi"

ganeshie8 (ganeshie8):

i don't see any pies in your options ?

OpenStudy (freckles):

maybe it meant everything that can be written in terms of pi not things that cannot be written in terms of pi

OpenStudy (anonymous):

i dont know?! you think if i grasped it properly i'd be asking basic trig questions on this website?

OpenStudy (freckles):

like for example arcsin(1)=pi/2 arcsin(1) can be written in terms of of pi some people would think an approximation to arcsin(1) would be ok if the teacher does not express she wants exact answer or the answer in terms of pi

OpenStudy (freckles):

if it was me I would probably have said instead express all answers in exact form

OpenStudy (freckles):

but then I guess you could say arcsin(1)=90 deg

OpenStudy (freckles):

and maybe that is why she said in terms of pi